OpenStudy (anonymous):

I need help finding the inverse of f(x) = 7 + sqrt(2 x - 5)

OpenStudy (anonymous):

so far I have \[y ^{2}-7^{2}=2x-5\]

OpenStudy (anonymous):

Let f(x) be y (for making it easy ) y - 7 =sqrt(2x-5) (y-7)^2 =2x - 5 \[\frac{(y-7)^2+5}{2} =x \] \[f^{-1}(x)= \frac{(x-7)^2+5 }{2}\]